What is the value of n for the reaction below?

NH4NO3(s) ⇌ N2O(g) + 2 H2O(g)

Kp is related to Kc by the equation Kp = Kc(RT)Δn.

A. -1

B. 0

C. -2

D. +2

E. +1
Answer #1

n = (sum of stoichiometric coefficients of products) - (sum of stoichiometric coefficients of reactants)

Here, N​​​​​​2​​​​​O and H​​​​2O are present on product side and their stoichiometric coefficients in the balanced reaction are 1 and 2 respectively.

Therefore, sum of stoichiometric coefficients of products = 1+2 = 3

Also, NH​​​​​4​NO​​3 is present on reactant side and its stoichiometric coefficient is 1.

Therefore, sum of stoichiometric coefficients of reactants = 1

Hence, n = 3-1 = +2.

So, D is the right answer.